40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 9 22:51:37 2013, Koji, Update, LSC, PRMI(sb) lock recovered 
    Reply  Mon Aug 12 10:44:22 2013, rana, Update, LSC, PRMI(sb) lock recovered 
    Reply  Mon Aug 12 14:05:34 2013, Jenne, Update, ASC, PRCL ASS software in place 
       Reply  Mon Aug 12 15:39:40 2013, Jenne, Update, ASC, PRCL ASS screens in place 
       Reply  Mon Aug 12 17:30:03 2013, Koji, Update, ASC, PRCL ASS software in place 
          Reply  Mon Aug 12 23:13:14 2013, Jenne, Update, ASC, PRCL ASS software in place 
Message ID: 8992     Entry time: Fri Aug 9 22:51:37 2013     Reply to this: 8994   8997
Author: Koji 
Type: Update 
Category: LSC 
Subject: PRMI(sb) lock recovered 

PRMI(sb) lock was recovered


PRMI lock

- Stared at the time series data of the REFL demod signals, and decided to use REFL165I&Q for the locking.

- Jiggled the demodulation phase of REFL165 and POP110. Changed the servo gains.

- Finally found a short lock. Further optimized the parameters.

- PRM ASC was turned on by giving the identity matrices for the input and output matrices.
  Now just hitting the up button is sufficient to engage the ASC servo.

- Under the presence of the ASC, the PRMI is indefinitely locked as before.

- Reacquisition is also instantaneous. (It acquires even if the ASC is left "on".)

- Actually the lock is somewhat robust even when the PRM ASC is not used.
  This is VERY GOOD as we can skip one of the steps necessary for the full lock.

  Although, the seismic on Friday night is very quiet.
  The spot motion at POP seems to be somewhat pitch/yaw mixed, in stead of previous "totally-dominated-by-yaw" situation.

- We are ready to implement ASS for PRM

Demod phase adjustment

- Shook PRM at 580Hz / 100cnt

- Swept the demod phase of REFL165 such that the PRM peak is minimized in the Q signal

- Open DTT. Measured transfer functions between REFL165I and the Q signals of each PD.

- Minimized the PRCL signal coupling in the signals.

- The resolution of the adjustment was ~1deg.

Locking test with PRM/BS

Tried the lock acquisition only with PRM and BS. (cf. http://nodus.ligo.caltech.edu:8080/40m/8816)

This just worked nicely.


Today's locking parameters:

PRMI(sb) lock:

MC Trans: 17500
POP110I (in lock): 150

PRCL Source: REFL165(I) 106deg / 45dB / Normalization SQRT(10 POP110I) / Input MTRX 1.0
PRCL Trigger: POP110I x 1.0 50up 25down
PRCL Servo: G=+3.5 Acq: FM4/FM5 Opr: FM2/FM3/FM6/FM7
PRCL Actuator: PRM +1.0

MICH Source: REFL165(Q) 106deg / 45dB / Normalization SQRT(0.1 POP110I) / Input MTRX 1.0
MICH Trigger: POP110I x 1.0 50up 25down
MICH Servo: G=-10 Acq: FM4/FM5 Opr: FM2/FM3/FM6
MICH Actuator: (ITMX -1.0 / ITMY +1.0) or (BS 0.5 / PRM -0.267)

Demod phases:

AS55 -17deg
REFL11 135deg
REFL33 -18deg
REFL55 120deg
REFL165 106deg

ELOG V3.1.3-